OpenStudy (anonymous):

If 11,000=8.2 and 13,000=6.9 what does 12,500=

OpenStudy (anonymous):

is there any formulas?

OpenStudy (anonymous):

4.7 maybe

OpenStudy (anonymous):

I was thinking IN between would be 7. Something

OpenStudy (anonymous):

hmm your right

OpenStudy (anonymous):

Trying to figure out a seeding population

OpenStudy (anonymous):

7.5 or 7.6 i think

OpenStudy (anonymous):

12000 is in between 7.6 and 7.5 just gotta figure out 12,500 for the exact 7.?

OpenStudy (anonymous):

8.2- 6.9 =1.3 drop in 2000 (13000-11000) 1.3/2000=x/500 ( 13000-12500)=500 x=7.225

OpenStudy (anonymous):

I belive that is it or close enough for me to figure up my soybean planting population to the acre thanks a lot Yall first time I used this Its just sometime my mind doesn't want to work right!!!!!